# Instructions for EMNLP 2023 Proceedings

### v1

@@ 5 Problems with the Minecraft Corpus
As a result, the net-action F1 metric, which evaluates a model’s action sequence based on whether the block placements match exactly in terms of block color and coordinates with the corresponding gold builder action, is often inappropriate. For instance, if the Builder puts down a block at one corner after receiving the instruction in a corner whereas NeBuLa chooses another corner, the metric would give NeBuLa zero credit whereas intuitively it still did the right thing. To summarize, the evaluation metric treats vague instructions as completely precise ones, and considers one instantiation of an instruction (i.e. the action sequence of Builder in the gold data) to be the only ground truth. Another related issue is highlighted in Figure 2 where the action sequence for the Architect’s instruction gets truncated by a question from the Builder “there?". In this case, for the aforementioned instruction, only the first three actions (place yellow -1 1 0, pick -1 1 0, place yellow -1 4 0) constitute the ground truth.
@@
To conclude, the underspecified instructions with multiple plausible instantiations, coupled with the strict nature of the metric, puts an upper bound on how much the net-action F1 score can improve on this dataset. More importantly, it doesn’t reveal what a model with a high F1 score actually does learn. We attempt to answer this in the next section.
6 Evaluating NeBuLa on Synthetic Dataset
@@
Given the issues associated with Minecraft Corpus and the evaluation metric, we test NeBuLa on simple scenarios using a more just metric. We begin by testing NeBuLa’s ability to construct simple shapes, such as, square, row, rectangle, tower, diagonal, diamond, cube of specific size and understand location (i.e. corner, centre, edge) and orientation descriptions (i.e. horizontal/vertical). We refer to all these shapes as level-1 structures. To do so, we construct a level-1 dataset of $1368$ instructions. Some of these instructions simply ask to construct a shape of specific size like “Build a $3\times 3$ red square.", while others are more detailed, for example, “Build a $3\times 3$ red horizontal square at the centre."
@@
For rows/diagonals/towers, we vary size from $3$ to $9$. For squares, the size varies from $3\times 3$ to $5\times 5$. For cubes, we only use $3\times 3\times 3$. For rectangles, we use sizes $m\times n$, where $m\neq n$, $m\times n<30$ and $84<=m<=$ 8. For diamonds, we use two variants to describe size “$m$ blocks on a side" and “axes $12m+$ 1 long", where $63<=m<=$ 6. We use orientation descriptions (i.e. horizontal/vertical) for squares, rectangles, and diamonds.
@@
shape corr# acc-shape acc-size loc-spec acc-loc or-spec acc-or
Tower 504 100% 100% 378 56.0%
Row 168 100% 100% 126 30.0%
diagonal 168 78.6% 95.0% 102 2.0%
rectangle 140 39.6% 12.0% 44 7.0% 31 100%
square 216 59.3% 96.0% 88 26.0% 75 81.0%
cube 24 58.3% 85.0% 8 37.0%
diamond 144 0% 0%
total 1368 73.0% 83.0% 746 38.0% 106 86%
Table 4: Evaluation of NeBuLa on shapes and basic locations. Shape accuracy gives how many of the shapes NeBuLa made were correct given the instruction. Additionally: acc-size–of those correct shapes how many were of the correct size; loc-spec–of the correct shapes how many had location specified; loc acc–of those correct shapes with specified locations how many were correctly located. We also tested rectangle and square for orientation (horizontal or vertical).
@@
Table 4 gives the result of NeBuLa on level-1 dataset. We don’t report color accuracy in the table, as NeBuLa always got the color correct. From the table, we can see that NeBuLa already has a decent command of basic shapes like towers, rows, and diagonals. However, it struggled with shapes like rectangle, square, cube, and diamond. It never correctly constructed diamonds, which might be because there were very few instances of diamonds in Minecraft corpus. For squares and rectangles which were correctly predicted, the model scored very high on orientation accuracy. However, the model has quite low location accuracy across all the correctly predicted shapes. The model rarely achieved an accuracy of above $50\%$, even with our relaxed evaluation method for locations.
@@
As a second step, we look at NeBuLa’s ability to understand location descriptions, in particular ones that are anaphorically specified. To do so, we start with an instantiation (randomly chosen from the set of correct instantiations) for the $1368$ instructions in level-1 dataset. So, for a level-1 instruction such as “Build a $3\times 3$ red square.", we have a $3\times 3$ red square already present in the grid. Now given a level-1 structure in the grid, we design level-2 instructions which require placing or removal of a specific color block. For place instructions, we use location descriptions like on top of, to the side of, touching, and not touching. So an example of level-2 place instruction is “place a blue block on top of that." where that refers to the level-1 structure in the grid. Similarly, for removal instructions, we have the simple instruction “remove a block" and more complex instructions including location descriptions like you just placed. We also have additional location descriptions for certain level-1 structures such as end for rows, diagonals; top, bottom for towers; corner for cube; centre for cube, odd-size squares and towers. An example of level-2 remove instruction is “remove the top block ."
@@
Similar to level-1, we evaluate NeBuLa on level-2 dataset by making use of binary functions like $is\_ontopof(b,C)$, $is\_touching(b,C)$ where $C$ is the level-1 structure already present in the grid and $b$ is the predicted block. As an example, for on top of, we check whether there is no block in C which is directly above the block b, and there is a block in C underneath block b.
@@
Table LABEL:tab:loc:eval shows that the model did quite well, with the exception of the instruction involving not touching as location description. Otherwise, the results indicate that NeBuLa has a good knowledge of basic anaphoric location descriptions.
@@
We then examined NeBuLa’s errors with on top of. We found that the failure cases mostly were a result of the model placing multiple blocks instead of just one on the given level-1 structure. That is, the model does not always understand a block as a single block. In light of these cases, when we check whether all the blocks in predicted $b$ are on top of $C$, the accuracy improves from $74.2\%$ to $97.2\%$. Thus, some of the difficulties NeBuLa had with instructions come from what might be a limited understanding of the semantics and pragmatics of indefinite and numerical noun phrases.
@@ 6.1 Finetuning NeBuLa on Shapes and Locations
shape tot# acc-shape acc-size loc-spec acc-loc or-spec acc-or
Tower 504 99.0% 100% 377 42.0%
Row 168 99.0% 100% 125 48.0%
diagonal 168 74.0% 80.0% 101 39.0%
rectangle 102 95.0% 49.0% 76 32.0% 65 100%
square 144 89.0% 100% 93 45.0% 86 100%
cube 24 100% 100% 18 66.0%
diamond 108 18.0% 0% 12 100%
total 1218 87.0% 90.0% 715 46.0% 163 100%
Table 6: Evaluation of NeBuLa after finetuning on shapes and basic locations. Shape accuracy (acc-shape), acc-size, location-spec, acc-loc, or-spec and acc-or as in Table 4
Our evaluation on level-1 and level-2 data shows that NeBuLa struggles with squares, rectangles, diamonds, and “not touching" place instructions. To tackle this, we used a subset of the two datasets to augment the training data for NeBuLa. From level-1 data, we took the following subset for training: squares of size $3\times 3$, diamonds of size $3$ (or axes $5$ spaces long), and rectangles of sizes $4\times 3$ and $5\times 4$. From level-2 data, we took those “touching/not touching” instances where the level-1 structure is square or rectangle. Out of total $363$ instances for touching/not touching, there were $109$ such instances. We then finetuned NeBuLa by combining the Minecraft training with this subset of level-1 and level-2 data. The rest of the level-1 and level-2 data was used for testing.
Table 6 shows NeBuLa’s performance on level-1 test set after finetuning. As before, we found that NeBuLa always got the color correct. From the table, we can see that the shape accuracy improved significantly for squares, rectangles, and diamonds in comparison with Table 4. Although the location accuracy is still low, it has improved in comparison with original NeBuLa. Interestingly, we also see that NeBuLa has perfect shape accuracy on cube although cube was not part of the training set. Finally, for correctly predicted shapes, NeBuLa achieved a perfect orientation accuracy.
Table LABEL:tab:loc:fine-tune shows the results on level-2 test set for NeBuLa after finetuning. Here also, we can see that NeBuLa’s accuracy remains very high on almost all of the simple instructions with the anaphoric location descriptions. Furthermore, its accuracy increased significantly for “not touching" instructions. This jump in accuracy is significant enough to conclude that NeBuLa has learned the concept of “contact", at least for our synthetic dataset. On the minecraft test set, we found that NeBuLa’s performance remained high with an average precision of $0.40$, recall of $0.414$ and a net action F1 of $0.391$. As we can see, these scores are at-par with the baseline NeBuLa.
@@ 7 Conclusions and Future Work
We have introduced NeBuLa, an LLM based action prediction model, for the Minecraft Dialogue Corpus. As a baseline, NeBuLa uses the entire Minecraft dialogue up to action $a_{n}$ to predict $a_{n}$. We showed that this baseline doubles the net action F1 scores of Jayannavar et al., (2020). We then showed that certain discourse structures provided necessary and sufficient information for inferring actions to the level of the baseline. We also analyzed NeBuLa’s errors on Minecraft corpus and provided additional finetuning to improve the model’s ability to interpret underspecified shape descriptions and anaphorically-specified locations using our synthetic dataset. This allowed us to analyze the shortcomings of the net-action F1 metric, and address them using a more realistic evaluation metric. Our evaluation metric captures the notion of relative location, but leaves exact locations typically underspecified, in accordance with our semantic intuitions. For future work, we plan to apply this metric (or a similar relative location metric) on the Minecraft corpus. Given the improvement in performance of NeBuLa after finetuning on our synthetic dataset, we hypothesize that in a more controlled collaborative task, with some pedagogical instructions to the Architect, NeBuLa could contribute as a useful interface for conversational robots that interact with humans.
@@
The MSDC contains a great deal of discourse information, including a full discourse structure analysis. We have only used some of this information. Potentially, we could leverage more information from this dataset to improve NeBuLa’s action prediction performance. We also need to extend our constraints to cover other frequent anaphoric location descriptions in addition to on top of X and to the side of X. Locutions like in front of/ behind, underneath, hanging off, next to $(X)$ all have underspecified parameters of either orientation, distance or direction that allow for several correct placements, once $X$ has been identified. We need to evaluate NeBuLa on these expressions as well. Finally, we need to reevaluate NeBuLa’s predictions as well as builder actions in the MDC with our more appropriate metric, which is suited to the underspecified shape and location descriptions used in the corpus.
@@
Our work here has been to improve the capacities of AI systems in interactive tasks where conversation can be used to optimize performance on collaborative actions. We see no direct ethical concerns that arise from this work. However, conversationally more capable robots, which could be one downstream application of this work, might require additional conversational strategies as constraints to ensure that participating humans retain the final say with regards to the actions in the collaborative tasks.
@@ 4 NVIDIA Volta V100
quantization for LLaMA3 4-bit NormalFloat
@@
Table 8 gives the hyperparameters used for finetuning NeBuLa along with the computing resources. We adapted the finetuning code from the following repository [^2] https://github.com/mlabonne/llm-course/blob/main/Fine_tune_Llama_2_in_Google_Colab.ipynb. We provide level-1 and level-2 synthetic data and NeBuLa trained on minecraft and the synthetic data here [^3] https://huggingface.co/akshay107/NeBuLa.

### v2

@@ 5 Problems with the MDC
As a result, the net-action F1 metric, which evaluates a model’s action sequence based on whether the block placements match exactly in terms of block color and coordinates with the corresponding gold builder action, is often inappropriate. For instance, if the Builder puts down a block at one corner after receiving the instruction in a corner whereas Nebula chooses another corner, the metric would give Nebula zero credit whereas intuitively it still executed the instruction correctly. To summarize, the net-action F1 evaluation metric treats vague instructions as completely precise ones, and considers one instantiation of an instruction (i.e. the action sequence of Builder in the gold data) to be the only ground truth. Another related issue is highlighted in Figure 2, where the action sequence for the Architect’s instruction gets truncated by a question from the Builder “there?”. In this case, for the aforementioned instruction, only the first three actions (place yellow -1 1 0, pick -1 1 0, place yellow -1 4 0) constitute the ground truth.
@@
Thus, the underspecified instructions with multiple plausible instantiations, coupled with the strict nature of the metric, puts an upper bound on how much the net-action F1 score can improve on this dataset. More importantly, it doesn’t reveal what a model with a high F1 score actually does learn. We attempt to answer this in the next section.
6 Evaluating Nebula on Synthetic Data
@@
Given the issues associated with MDC and the evaluation metric, we test baseline Nebula (i.e., the one trained on entire conversation history) on simple scenarios using a more just metric. To do so, we construct synthetic datasets at two different levels. The goal of these datasets is to test what basic shapes and location descriptors Nebula learns after being trained on MDC. For the first level, we test Nebula’s ability to construct simple shapes, such as, square, row, rectangle, tower, diagonal, diamond, cube of specific size and understand location (i.e. corner, centre, edge) and orientation descriptions (i.e. horizontal/vertical). We refer to all these shapes as level-1 structures. The resulting dataset, referred to as level-1 dataset, consists of $1368$ instructions. Some of these instructions simply ask to construct a shape of specific size like “Build a $3\times 3$ red square.”, while others are more detailed, for example, “Build a $3\times 3$ red horizontal square at the centre.”
@@
For rows/diagonals/towers, we vary size from $3$ to $9$. For squares, the size varies from $3\times 3$ to $5\times 5$. For cubes, we only use $3\times 3\times 3$. For rectangles, we use sizes $m\times n$, where $m\neq n$, $m\times n<30$ and $84<=m<=$ 8. For diamonds, we use two variants to describe size “$m$ blocks on a side” and “axes $12m+$ 1 long”, where $63<=m<=$ 6. We use orientation descriptions (i.e. horizontal/vertical) for squares, rectangles, and diamonds.
Shape Total # ShapeAccSizeAccLoc-spec LocAccOrient-spec OrientAcc%
BL FT BL FT BL FT BL FT BL FT BL FT BL FT
Tower 504 504 100 99.0 100 100 378 377 56.0 42.0
Row 168 168 100 99.0 100 100 126 125 30.0 48.0
Diagonal 168 168 78.6 74.0 95.0 80.0 102 101 2.0 39.0
Rectangle 140 102 39.6 95.0 12.0 49.0 44 76 7.0 32.0 31 65 100 100
Square 216 144 59.3 89.0 96.0 100 88 93 26.0 45.0 75 86 81.0 100
Cube 24 24 58.3 100 85.0 100 8 18 37.0 66.0
Diamond 144 108 0 18.0 0 0 12 100
Total 1368 1218 73.0 87.0 83.0 90.0 746 715 38.0 46.0 106 163 86.0 200
Table 4: Evaluation of baseline Nebula (BL) and Nebula finetuned further on a part of synthetic data (FT), on shapes and basic locations. ShapeAccgives percentage of cases where the given shape was correct. Additionally, SizeAccdenotes, for the correct shapes, percentage of cases where it is of the correct size; Loc-spec denotes, for the correct shapes, how many have location specified; LocAccdenotes location accuracy for such cases. We also test rectangle and square for orientation (horizontal or vertical). Orient-spec denotes, for the correct shapes, the number of cases where orientation is specified; and OrientAccdenotes the orientation accuracy for the same.
Instruction Total # Accuracy(%)
BL FT BL FT
Overall 1368 1259 80.4 89.6
Place…
on top of 178 178 74.2 79.7
to the side of 154 154 98.1 87.7
touching 176 120 99.4 93.3
not touching 187 134 7.5 97.8
Place Overall 695 586 67.9 88.7
Remove…
any block 234 234 95.3 94.4
block just placed 216 216 95.3 84.7
top block 44 44 100 100
bottom block 65 65 100 100
centre block 56 56 60.7 66.1
corner block 2 2 100 100
end block 56 56 96.4 100
Remove Overall 673 673 93.3 90.3
Table 5: Evaluation of baseline Nebula (BL) and Nebula finetuned further on a part of synthetic data (FT), on location descriptors for place and remove instructions. The FT model performs considerably better for “not touching” place instructions while remaining at-par for other instruction types.
@@
Table 4 gives the result of baseline Nebula (refer to BL in the table) on level-1 dataset. We don’t report color accuracy, as Nebula always gets the color correct. From the table, we can see that Nebula already has a decent command of basic shapes like towers, rows, and diagonals. However, it struggles with shapes like rectangle, square, cube, and diamond. It never correctly constructs diamonds, which might be because there are very few instances of diamonds in MDC. For squares and rectangles which are correctly predicted, the model scores very high on orientation accuracy. However, the model has quite low location accuracy across all the correctly predicted shapes. The model rarely achieves an accuracy of above $50\%$, even with our relaxed evaluation method for locations.
We also evaluate the performance of Neural Builder Jayannavar et al., (2020) on our level-1 dataset. The results are shown in Table 7 of the Appendix. Apart from rows and towers, Neural Builder model fails to build any shape correctly.
@@
As a second step, we look at Nebula’s ability to understand location descriptions, in particular ones that are anaphorically specified. To do so, we start with an instantiation (randomly chosen from the set of correct instantiations) for the $1368$ instructions in level-1 dataset. So, for a level-1 instruction such as “Build a $3\times 3$ red square.”, we have a $3\times 3$ red square already present in the grid. Now given a level-1 structure in the grid, we design level-2 instructions which require placing or removal of a specific color block. For place instructions, we use location descriptions like on top of, to the side of, touching, and not touching. So an example of level-2 place instruction is “place a blue block on top of that.” where that refers to the level-1 structure in the grid. Similarly, for removal instructions, we have the simple instruction “remove a block” and more complex instructions including location descriptions like you just placed. We also have additional location descriptions for certain level-1 structures such as end for rows, diagonals; top, bottom for towers; corner for cube; centre for cube, odd-size squares and towers. An example of level-2 remove instruction is “remove the top block.” Both level-1 and level-2 datasets were generated in an automated fashion.
@@
Similar to level-1, we evaluate Nebula on level-2 dataset by making use of binary functions like $is\_ontopof(b,C)$, $is\_touching(b,C)$ where $C$ is the level-1 structure already present in the grid and $b$ is the predicted block. For example, binary function for on top of checks whether there is no block in C which is directly above the block b, and there is a block in C underneath block b.
@@
Table 5 shows that baseline Nebula performs quite well (refer to BL in the table), with the exception of the instruction involving not touching as location description. These results indicate that Nebula has a good knowledge of basic anaphoric location descriptions. Similar to level-1 dataset, we look at the performance of Neural Builder Jayannavar et al., (2020) on the level-2 dataset. Table 8 in the Appendix shows that the performance of the Neural Builder is considerably worse than Nebula for all the location descriptors.
@@
We then examine Nebula’s errors with on top of. We find that the failure cases mostly were a result of the model placing multiple blocks instead of just one on the given level-1 structure. That is, the model does not always understand a block as a single block. In light of these cases, when we check whether all the blocks in predicted $b$ are on top of $C$, the accuracy improves from $74.2\%$ to $97.2\%$. Thus, some of the difficulties Nebula had with instructions come from what might be a limited understanding of the semantics and pragmatics of indefinite and numerical noun phrases. We provide visual comparisons of Neural Builder and baseline Nebula on level-1 and level-2 dataset examples in Tables 9 and 10 of the Appendix.
@@ 6.1 Finetuning Nebula on Shapes and Locations
Our evaluation on level-1 and level-2 data shows that Nebula struggles with squares, rectangles, diamonds, and “not touching” place instructions. To tackle this, we use a subset of the two datasets to augment the training data for Nebula. From level-1 data, we take the following subset for training: squares of size $3\times 3$, diamonds of size $3$ (or axes $5$ spaces long), and rectangles of sizes $4\times 3$ and $5\times 4$. From level-2 data, we take those “touching/not touching” instances where the level-1 structure is square or rectangle. Out of total $363$ instances for touching/not touching, there are $109$ such instances. We then finetune Nebula by combining the Minecraft training with this subset of level-1 and level-2 data. The rest of the level-1 and level-2 data is used for testing.
Table 4 shows Nebula’s performance on the level-1 test set after finetuning (refer to FT in the table). As before, we find that Nebula always got the color correct. From the table, we can see that the shape accuracy improves significantly for squares, rectangles, and diamonds in comparison to baseline Nebula (BL in the table). Although the location accuracy is still low, it improves in comparison with baseline Nebula. Interestingly, we also see that Nebula has perfect shape accuracy on cube, although cube is not part of the training set. Finally, for correctly predicted shapes, Nebula achieves a perfect orientation accuracy.
Table 5 shows the results on the level-2 test set for Nebula after finetuning. Here also, we can see that Nebula’s accuracy remains very high on almost all of the simple instructions with the anaphoric location descriptions. Furthermore, its accuracy increases drastically for “not touching” instructions. This jump in accuracy is significant enough to conclude that Nebula has learned the concept of “contact”, at least for our synthetic dataset. On the minecraft test set, we find that Nebula’s performance remains high with a net action F1 of $0.391$. As we can see, these scores are at-par with the baseline Nebula (refer to llama-3-8b in Table 2).
@@ 7 Conclusions and Future Work
We introduce Nebula, an LLM based action prediction model, for the Minecraft Dialogue Corpus. As a baseline, Nebula uses the entire Minecraft dialogue up to action $a_{n}$ to predict $a_{n}$. We show that this baseline doubles the net action F1 scores of Jayannavar et al., (2020). We then show that certain discourse structures provide necessary and sufficient information for inferring actions at-par with the baseline setting. We also analyze Nebula’s errors on MDC and provide additional finetuning to improve the model’s ability to interpret underspecified shape descriptions and anaphorically-specified locations using our synthetic dataset. This allows us to analyze the shortcomings of the net-action F1 metric, and address them using a more realistic evaluation metric. Our evaluation metric captures the notion of relative location, but leaves exact locations typically underspecified, in accordance with our semantic intuitions. For future work, we plan to explore metrics similar to our relative location metric that can be applied more generally, including on the MDC. Given the improvement in performance of Nebula after finetuning on our synthetic dataset, we hypothesize that in a more controlled collaborative task, with some pedagogical instructions to the Architect, Nebula could contribute as a useful interface for conversational robots that interact with humans.
@@
The MSDC contains a great deal of discourse information, including a full discourse structure analysis. We only use some of this information. Potentially, we could leverage more information from this dataset to improve Nebula’s action prediction performance. We also need to extend our constraints to cover other frequent anaphoric location descriptions in addition to on top of X and to the side of X. Locutions like in front of/ behind, underneath, hanging off, next to $(X)$ all have underspecified parameters of either orientation, distance or direction that allow for several correct placements, once $X$ has been identified. We need to evaluate Nebula on these expressions as well. Finally, we need to reevaluate Nebula’s predictions as well as the original builder actions in the MDC with our more appropriate metric, which is suited to the underspecified shape and location descriptions used in the corpus.
@@
Our work here has been to improve the capacities of AI systems in interactive tasks where conversation can be used to optimize performance on collaborative actions. We see no direct ethical concerns that arise from this work. Though conversationally more capable robots, which could be one downstream application of this work, might require additional conversational strategies as constraints to ensure that participating humans retain the final say with regards to the actions in the collaborative tasks.
Acknowledgements
For financial support, we thank the National Interdisciplinary Artificial Intelligence Institute ANITI (Artificial and Natural Intelligence Toulouse Institute), funded by the French ‘Investing for the Future–PIA3’ program under the Grant agreement ANR-19-PI3A-000. We also thank the projects COCOBOTS (ANR-21-FAI2-0005) and DISCUTER (ANR-21-ASIA-0005), and the COCOPIL “Graine” project of the Région Occitanie of France. This project has also been funded by the France 2030 program and is funded by the European Union - Next Generation EU as part of the France Relance. This work was granted access to the HPC resources of CALMIP supercomputing center under the allocation 2016-P23060. We would also like to thank Bastien Navarri for designing the visualization software used in Tables 9 and 10.
@@ 4 NVIDIA Volta V100
quantization for Llama-3 4-bit NormalFloat
@@
Table 6 gives the hyperparameters used for finetuning Nebula along with the computing resources. We adapt the finetuning code from the following repository [^2] https://github.com/mlabonne/llm-course/blob/main/Fine_tune_Llama_2_in_Google_Colab.ipynb.
Shape Total # ShapeAccSizeAccLoc-spec LocAccOrient-spec OrientAcc%
Tower 504 75.6 20.5 303 0 0 0
Row 168 97.6 16.6 126 32.5 0 0
Diagonal 168 0 0 0 0 0 0
Rectangle 140 0 0 0 0 0 0
Square 216 0 0 0 0 0 0
Cube 24 0 0 0 0 0 0
Diamond 144 0 0 0 0 0 0
Total 1368 39.8 19.4 429 9.5 0 0
Table 7: Evaluation of Neural Builder Jayannavar et al., (2020) on shapes and basic locations. ShapeAccgives percentage of cases where the given shape was correct. Additionally, SizeAccdenotes, for the correct shapes, percentage of cases where it was of the correct size; Loc-spec denotes, for the correct shapes, how many had location specified; LocAccdenotes location accuracy for such cases. We also tested rectangle and square for orientation (horizontal or vertical). Orient-spec denotes, for the correct shapes, the number of cases where orientation was specified; and OrientAccdenotes the orientation accuracy for the same.
Instruction Total # Accuracy(%)
Overall 1368 50.9
Place…
on top of 178 60.1
to the side of 154 66.2
touching 176 98.3
not touching 187 0.5
Place Overall 695 55.1
Remove…
any block 234 88.5
block just placed 216 14.8
top block 44 2.3
bottom block 65 35.4
centre block 56 16.1
corner block 2 0
end block 56 73.2
Remove Overall 673 46.5
Table 8: Evaluation of Neural Builder Jayannavar et al., (2020) on location descriptors for place and remove instructions.
Lvl-1 Instruction Neural Builder Nebula Build a purple tower of size 3 at the centre. [Uncaptioned image] [Uncaptioned image] Build a 3x3 purple square at the centre. [Uncaptioned image] [Uncaptioned image] Build a diagonal of 3 purple blocks at an edge. [Uncaptioned image] [Uncaptioned image] Build a 6x3 green rectangle at the centre. [Uncaptioned image] [Uncaptioned image]
Table 9: Comparison of Neural Builder Jayannavar et al., (2020) and baseline Nebula on level-1 dataset.
World State + Lvl-2 Instruction Neural Builder Nebula [Uncaptioned image] OK now remove the centre block. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] OK now place a yellow block to the side of that. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] OK now remove the top block. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] OK now remove the bottom block. [Uncaptioned image] [Uncaptioned image]
Table 10: Comparison of Neural Builder Jayannavar et al., (2020) and baseline Nebula on level-2 dataset.